OpenStudy (anonymous):

The point (1.5, -1.5) is the midpoint of (-1,5) and what other point?

OpenStudy (muhammad_nauman_umair):

let p(x,y) be the other point (1.5,-1.5)=((-1+x)/2,(5+y)/2) 2*1.5=-1+x x=4 2*-1.5=5+y y=-8